OpenStudy (anonymous):

|| HELP PLEASE!!!|| What is the solution to the system of equations? 5x - 3y = -9, 2x - 5y = 4

OpenStudy (solomonzelman):

\(\large\color{slate}{ \left[\begin{matrix}~5~ & ~-3~ & |&~-9 \\ ~2~ & ~-5~ & |&~4 \end{matrix}\right] }\) don't know why matrix appears to be good to me, but I would consider that option.

OpenStudy (anonymous):

i got y = -2, x= -3 but idk if it's right.

OpenStudy (texaschic101):

5x - 3y = -9 --->(2) 2x - 5y = 4 --->(-5) ----------------- 10x - 6y = -18 (result of multiplying by 2) -10x + 25y = -20 (result of multiplying by -5) -----------------add 19y = -38 y = -2 5x - 3y = -9 5x - 3(-2) = -9 5x + 6 = -9 5x = -9 - 6 5x = -15 x = -3 check.. 2x - 5y = 4 2(-3) - 5(-2) = 4 -6 + 10 = 4 4 = 4 (correct) so x = -3 and y = -2

OpenStudy (texaschic101):

you got what I got :)

OpenStudy (texaschic101):

you can always check your answer by subbing it back into either equation to see of they come out equal...if they do, then it is correct

OpenStudy (anonymous):

Thank you so much! @texaschic101 & @SolomonZelman

OpenStudy (texaschic101):

sure thing :) I didn't really do anything...just checked your work and found it to be correct...you did the work

OpenStudy (anonymous):

Do you think you could help me with another question? I'm not sure how to do it at all and I just need to be walked through it. Please? @texaschic101

OpenStudy (texaschic101):

whats the problem ?

OpenStudy (anonymous):

:\\

OpenStudy (texaschic101):

inconsistent means there is no solution...it means that the lines are parallel...they have the same slopes but different y intercepts. In y = mx + b form, the slope is in the m position and the y intercept is in the b position. This one is confusing me a little bit...you might want to ask Solomon